use the method of disks to determine the volume of the solid by rotating the region bounded by y=10-2x,y=x+1 and y=7 about the

a. line x=8 b. line x=1 c. line x=-4


1
Expert's answer
2022-05-24T11:40:31-0400
"y=10-2x=>x=-0.5y+5"

"y=x+1=>x=y-1"


"y-1=-0.5y+5"

"y=4"

a.


"V=\displaystyle\int_{4}^{7}\pi(8+0.5y-5)^2dy""-\displaystyle\int_{4}^{7}\pi(8-y+1)^2dy"

"=\pi[-0.25y^3+10.5y^2-72y]\begin{matrix}
 7\\
4
\end{matrix}="

"=\dfrac{243\pi}{4}({units}^3)"

b.


"V=\displaystyle\int_{4}^{7}\pi(y-1-1)^2dy""-\displaystyle\int_{4}^{7}\pi(-0.5y+54)^2dy"

"=\pi[0.25y^3-12y]\begin{matrix}
 7\\
4
\end{matrix}="

"=\dfrac{135\pi}{4}({units}^3)"

c.


"V=\displaystyle\int_{4}^{7}\pi(y-1+4)^2dy""-\displaystyle\int_{4}^{7}\pi(-0.5y+5-1)^2dy"

"=\pi[0.25y^3+7.5y^2-72y]\begin{matrix}
 7\\
4
\end{matrix}="

"=\dfrac{405\pi}{4}({units}^3)"